# Revisiting MAE pre-training for 3D medical image segmentation

### v1

@@ 3 Revisiting 3D MAEs
Table 2: Fine-tuning matters. We compare various combinations of weight transfer and fine-tuning schedules. Transfer: [Uncaptioned image] Transfer all, [Uncaptioned image] Transfer encoder only. Warm-Up and Training: [Uncaptioned image] Only decoder weights trained, [Uncaptioned image] Encoder and decoder weights trained. $\ddagger$: nnU-Net default (Dynamic)
@@ 4 Results and Discussion
We compare our final models S3D-B and S3D-L against VoCo (Wu et al., 2024), VolumeFusion (VF) by Wang et al. (2023), as well as MG Models Genesis (Zhou et al., 2021). The baselines are pre-trained using the same framework on the same data with the same backbone and the same hyperparameters - where possible - and are scaled to fully utilize an A100 40GB GPU, optimized for 250k steps. We provide explicit baseline method and configuration details in Appendix A. Moreover, we compare against two from-scratch baselines. The first, ’No (Dyn.)’, represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings. The second from-scratch baseline, ’No (Fixed)’, is an nnU-Net training with the same plans and preprocessing as defined by our pre-training. The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Table 3. Additionally, ranking stability of the methods is evaluated through bootstrapping and is provided in Fig. 3.
@@ 4.1 Observations
Across all tested datasets, SSL pre-trained methods demonstrate improved downstream segmentation performance. Comparing our S3D-B method to the most similar from-scratch baseline ’No (Fixed)’, we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points. This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and a state-of-the-art architecture.
@@
Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG, S3D-B, and S3D-L) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training. Given the age of ’Models Genesis’ - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF. We attribute this to a combination of two factors: 1. Models Genesis was originally published and trained on an outdated 3D-UNet (Çiçek et al., 2016) and outside of the powerful nnU-Net framework (Isensee et al., 2021). This highlights the importance of avoiding Pitfall 2: Training on a state-of-the-art backbone. 2. VoCo and VF were introduced in conjunction with architectures they were optimized for. By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture-pretraining combination may be suboptimal for the new CNN backbone.
S3D-B ranks first with respect to DSC and Normalized Surface Distance (NSD), while S3D-L and MG share the second place. Although S3D-L and MG are very close in Average DSC, S3D-L consistently achieves lower ranks across all datasets. Moreover, according to the bootstrapped aggregated rank (Fig. 3), this superior ranking is reliably higher than that of MG. These results indicate the overall efficacy of our pre-training method compared to currently established methods for CNN pre-training.
SSL Method No (Dyn.) No (Fixed) VoCo VF MG S3D-B S3D-L
@@
MS FLAIR (D1) 57.81 59.82 59.70 59.29 58.64 60.35 59.85
Brain Mets (D2) 63.66 56.53 56.25 61.01 65.39 65.24 64.81
Hippocampus (D3) 89.18 89.24 88.78 89.03 89.38 89.60 89.34
Atlas22 (D4) 63.28 65.52 62.97 65.76 65.93 66.95 64.58
CrossModa (D5) 85.64 83.44 83.07 84.24 83.91 84.08 84.02
Cosmos22 (D6) 60.28 78.17 77.40 80.09 79.67 80.00 80.01
ISLES22 (D7) 77.94 79.44 78.14 78.96 78.85 79.70 79.89
Hanseg (D8) 59.00 61.85 57.47 61.49 62.52 62.11 61.93
HNTS-MRG24 (D9) 66.73 65.90 67.65 63.34 68.00 68.62 67.94
BRATS24 Africa (D10) 93.07 92.51 91.97 92.16 92.36 92.19 92.90
T2 Aneurysms (D11) 46.76 41.97 40.16 44.96 45.48 47.26 44.15
Avg. DSC 69.40 70.40 69.41 70.94 71.83 72.37 71.77
Avg. Rank 4.64 4.55 6.27 4.36 3.18 2.00 3.00
@@
MS FLAIR (D1) 78.77 80.16 79.70 79.57 79.16 80.03 80.40
Brain Mets (D2) 80.72 76.72 72.77 79.20 81.51 82.53 82.32
Hippocampus (D3) 99.46 99.42 99.43 99.46 99.39 99.46 99.44
Atlas22 (D4) 70.52 73.77 70.15 73.67 74.22 75.35 73.45
CrossModa (D5) 99.85 99.76 99.72 99.78 99.74 99.81 99.80
Cosmos22 (D6) 72.60 96.47 94.48 96.89 96.95 97.45 96.75
ISLES22 (D7) 88.55 90.45 89.39 90.28 89.59 90.59 90.72
Hanseg (D8) 82.20 85.94 80.44 85.29 85.94 85.80 86.20
HNTS-MRG24 (D9) 71.83 71.26 73.47 67.88 73.22 74.07 73.17
BRATS24 Africa (D10) 95.66 95.36 94.95 94.94 95.33 95.06 95.72
T2 Aneurysms (D11) 62.24 55.56 51.79 58.97 59.38 61.18 57.07
Avg. NSD 82.04 84.08 82.39 84.17 84.95 85.58 85.00
Avg. NSD Rank 4.27 4.27 5.82 4.64 4.00 2.18 2.82
Table 3: S3D out-performs all baselines: Mean DSC and NSD results across the test datasets, representing a broad selection of brain MR tasks, are presented. ’No Fixed’ represents a from-scratch baseline sharing the same architecture, preprocessing and downstream training steps as all SSL methods. ’No Dyn.’ represents the original nnU-Net adapted to each downstream dataset individually.
Figure 3: Dataset rank stability In addition to absolute mean performance, we report the ranking stability of the methods through bootstrapping for all test datasets as well as the aggregated rank across all datasets.
@@
Comparing the ’No (Dyn.)’ and ’No (Fixed)’ configurations, both trained from scratch, reveals that selecting the appropriate configuration for each dataset can significantly influence performance. For instance, on datasets D2 and D11, the dynamic configuration outperforms the fixed by +7 and +5 DSC points, respectively, while for D6, the fixed configuration yields results +18 DSC points higher. In the majority of datasets where the fixed configuration underperforms relative to the dynamic nnU-Net, pre-training helps to recover performance. However, in some cases, such as with D5, the dynamic default nnU-Net still proves superior.
@@ 4.2 Ablation experiments
Results are presented in Table 4. It is evident that our pre-trained S3D-B model leads to better downstream performance compared to training from scratch in this setting. With just 40 trained images, the fine-tuned model nearly matches the performance of the from-scratch model trained on the full dataset. Future research could explore whether optimizing the training duration or learning rate schedule could prevent the pre-trained network from overfitting during fine-tuning on such a limited number of training images.
@@
To assess the generalization capability of the proposed pre-training method, we tested two scenarios. First, we evaluated fine-tuning our method on an unseen modality using the TOF Angiography Aneurysms dataset (D12). As shown in Table 7, without pretraining, the fixed configuration suffers a performance drop of 20 Dice points. We attribute this to the significant difference in median spacing for the downstream task ($[0.50,0.43,0.43]$ mm), which has a higher resolution than the fixed target spacing of $[1,1,1]$ mm used in the pre-training experiments. This lower resolution likely increases the difficulty of segmenting small aneurysms. Despite this decrease, pre-training mitigates some of this degradation and proves highly beneficial compared to training the same configuration from scratch. Interestingly, ModelGenesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences. Second, we fine-tuned on the D2 dataset using only the T1 contrast-enhanced (T1ce) sequence and applied these models directly to D13 without any additional fine-tuning. While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that MAE pre-training improves generalization across different centers, with our S3D model yielding the best performance.
@@
Previous studies have demonstrated the positive impact of extended training schedules on the quality of learned representations for downstream tasks (He et al., 2022; Feichtenhofer et al., 2022). To explore the relevance of this factor in the 3D medical domain, we conducted a similar experiment, evaluating training durations ranging from 62.5k to 1M steps. Our results indicate that the benefits of longer training schedules begin to degrade after 250k steps, as evident in Table 5. This could explain why our scaled model did not achieve further performance improvements.
@@
Initializing from pre-trained weights has the potential to reduce the computational resources needed for the network to adapt to new tasks. To assess this, we tested different pre-training durations on our development datasets. While maintaining 12.5k iterations for both the decoder and full network warm-up phases, we experimented with varying subsequent training lengths. As shown in Table 8, adding just 12.5k additional iterations (37.5k total) already outperforms training from scratch. However, achieving optimal performance still requires completing the full fine-tuning schedule.
@@
In many medical examinations, it is common to perform multiple scans, as clinicians often require images with different characteristics for accurate decision-making. Consequently, some datasets, like D2, D8, and D10, contain multiple input modalities. While pre-training may involve all modalities, we only feed one modality at a time into the network since not all patients have scans in every modality. This raises the question of how to handle datasets with multiple registered images. To address this, we conducted a 5-fold cross-validation on the D2 development dataset. We evaluated the replication of each input modality along with random initialization of the input stem weights. Additionally, we tested freezing the stem weights during the decoder warm-up phase. As shown in Table 6, the most stable and consistently effective approach was replicating the pre-trained stem and keeping it frozen during the decoder’s warm-up period.
@@ 5 Conclusion
This work is the first to demonstrate the potential of properly configured MAEs in 3D medical image segmentation. By overcoming key pitfalls in previous research, such as limited dataset sizes, outdated architectures, and insufficient evaluation, we show a consistent performance improvement over previous SSL methods. Notably, for the first time, we achieve consistent gains over the dynamic, dataset-adaptive nnU-Net baseline, validated across a large and diverse set of development and testing datasets. While our findings are promising, several avenues remain open for future exploration. Notably, increasing training time and batch size did not lead to performance gains, but the question remains of whether scaling the pre-training dataset size or the model parameters could unlock new potential. Furthermore, the intensity shifts employed by ModelGenesis SSL task hint at intriguing possibilities for improving generalization across unseen MRI modalities, which needs to be further explored for MAEs. Lastly, a data-centric approach to curating the most relevant data for SSL represents an exciting frontier for future research. Raw clinical datasets often contain images not intended for diagnostic purposes, such as those used for scanner calibration, which can dilute the effectiveness of pre-training. While we applied basic filtering to exclude low-quality data, more sophisticated filtering techniques could significantly enhance the quality of the pre-training process.
This work follows the spirit of prior studies like nnU-Net Isensee et al. (2021), or XXX, showing that a robust development strategy, informed model configuration, and rigorous validation lead to true and sustainable performance gains, contrasting the current hype for employing and modifying the latest network architecture. With our dynamic open-source framework, we hope to contribute to a cultural shift in the community towards validation-driven development enabling true scientific progress.
@@ Appendix A Method configuration
Across all baseline methods, we utilize a common set of hyperparameters. For all baseline methods we utilize the same pre-training dataset with the same image preprocessing. Moreover we use the same amount of pre-training steps (250k) as for our S3D-B method and the same fine-tuning scheme, as highlighted in Table 8. Aside from this, we employ the SGD optimizer with LR 1e-2 with a PolyLR schedule, momentum 0.99 and weight decay 3e-5 across all pre-training experiments, as they showed to be highly robust and reliable in the supervised medical image segmentation setting using CNNs (Isensee et al., 2021). Moreover, we denote that all these methods have their backbones replaced with a ResEnc U-Net to minimize confounding effects of different architectures.
@@ A.1 Models Genesis
Models genesis (Zhou et al., 2021) pre-text task is centered around restoring original patches from transformed versions. The transformed version is achieved by applying four different transformations in various combinations, with the following transformations: a composition of four separate pre-training schemes: (i) Non-linear intensity transformation:Alters the intensity distribution while preserving the anatomy, focusing on learning the appearance of organs. (ii) Out-painting:Removes part of the image and requires the model to extrapolate from the remaining image, forcing it to learn the global structure of the organs. (iii) In-painting:Masks a part of the image, and the model learns to restore the missing parts, focusing on local continuity and context. After having transformed the original image through these augmentations, the model is trained to recover the original image through a convolutional encoder-decoder architecture. This approach consolidates different tasks (appearance, texture, and context learning) into one unified image restoration task, making the model more robust and generalizable.
@@
The entire set of hyperparameters of Models Genesis are contained within the data-augmentation. This allows us to transfer this transformation pipeline, as provided in the official https://github.com/MrGiovanni/ModelsGenesis without any changes to the hyperparameters.

### v2

@@ 4 Revisiting 3D MAEs
Fine-tuning strategy
Given a pre-trained model, a crucial question arises: Which weights to transfer and how to schedule the fine-tuning? We investigated various different schedules, with all of themimplemented in the nnU-Net framework [18]. Regarding weight transfer we investigated transferring (i) both the encoder and decoder [Uncaptioned image], or (ii) only the encoder [Uncaptioned image] with a randomly initialized decoder. Regarding the fine-tuning schedule, we investigate whether to use a learning rate warm-up of 12.5k steps, ramping the learning rate up to the maximum LR. When only transferring the encoder, an additional warm-up of only the decoder is investigated to adapt the randomly initialized decoder to the pre-trained encoder. In some configurations, this results in two learning rate warm-ups of 12.5k steps each. Additionally, we investigate whether to keep the encoder frozen [Uncaptioned image] for the entire fine-tuning process, or to fine-tuning encoder weights as well. Lastly, we investigate whether to decrease the peak learning rate to $31e-$ 3, $41e-$ 4 or to keep it at the default of $21e-$ 2.
@@ 5 Results and Discussion
We compare our final model S3D against Volume Contrastive (VoCo) [42], VolumeFusion (VF) by Wang et al. [39], as well as Models Genesis (MG) [44]. The baselines are pre-trained for 250k steps using the same framework on the same data with the same backbone and the same hyperparameters (where possible) but scaled to fully utilize an A100 40GB GPU. We provide explicit baseline methods and configuration details in Appendix A. Moreover, we compare against two strong from-scratch baselines. The first, No (Dyn.), represents a non-pretrained (i.e., from scratch) default nnU-Net 3D fullres architecture that was planned and preprocessed on each downstream dataset individually, potentially resulting in different architectures, data preprocessing, and spacings. The second from-scratch baseline, No (Fixed), is a nnU-Net training with the same plans and pre-processing as defined by our pre-training. The dataset-wise mean DSC and mean NSD values across our test dataset suite are provided in Tab. 4. Additionally, the ranking stability of the methods is evaluated through bootstrapping and is provided in Fig. 2.
@@ 5.1 Observations
Across all tested datasets, SSL pre-trained methods demonstrate improved downstream segmentation performance. Comparing our S3D method to the most similar from-scratch baseline No (Fixed), we observe higher DSC scores in 10 out of 11 test datasets, with an average increase of +2 DSC points and +1.6 NSD points. This improvement is not limited to our method; MG and VF also achieve higher performance than the baseline, indicating the utility of SSL methods when applied to sufficient data and state-of-the-art architecture.
@@
Throughout our test dataset pool, SSL schemes using the masked image modeling paradigm (MG and S3D) consistently rank higher than the contrastive VoCo or the pseudo-segmentation-based VolumeFusion pre-training method for CNN pre-training. Given the age of Models Genesis - published in 2019 - it is surprising to see it outperform the more recent VoCo or VF. We attribute this to a combination of two factors: 1. Models Genesis was originally published and trained on an outdated 3D-UNet [6] and outside of the powerful nnU-Net framework [18]. This highlights the importance of avoiding Pitfall 2: Training on a state-of-the-art backbone. 2. VoCo and VF were introduced in conjunction with architectures they were optimized for. By transferring them to a CNN setting, hyperparameters chosen to optimize the method for their original architecture pre-training combination may be suboptimal for the new CNN backbone.
Table 4: S3D out-performs all baselines: The upper table presents the DSC results, while the lower table displays the NSD results across the test datasets, covering a diverse range of brain MRI tasks. No Fixed represents a from-scratch baseline sharing the same architecture, preprocessing and downstream training steps as all SSL methods. No Dyn. represents the original nnU-Net adapted to each downstream dataset individually.
SSL Method No (Dyn.) No (Fix.) VoCo VF MG S3D
@@
MS FLAIR (D1) 57.81 59.82 59.70 59.29 58.64 60.35
Brain Mets (D2) 63.66 56.53 56.25 61.01 65.39 65.24
Hippocampus (D3) 89.18 89.24 88.78 89.03 89.38 89.60
Atlas22 (D4) 63.28 65.52 62.97 65.76 65.93 66.95
CrossModa (D5) 85.64 83.44 83.07 84.24 83.91 84.08
Cosmos22 (D6) 60.28 78.17 77.40 80.09 79.67 80.00
ISLES22 (D7) 77.94 79.44 78.14 78.96 78.85 79.70
Hanseg (D8) 59.00 61.85 57.47 61.49 62.52 62.11
HNTS-MRG24 (D9) 66.73 65.90 67.65 63.34 68.00 68.62
BRATS24 Africa (D10) 93.07 92.51 91.97 92.16 92.36 92.19
T2 Aneurysms (D11) 46.76 41.97 40.16 44.96 45.48 47.26
Avg. DSC 69.40 70.40 69.41 70.94 71.83 72.37
Avg. Rank 4.64 4.55 6.27 4.36 3.18 2.00
@@
MS FLAIR (D1) 78.77 80.16 79.70 79.57 79.16 80.03
Brain Mets (D2) 80.72 76.72 72.77 79.20 81.51 82.53
Hippocampus (D3) 99.46 99.42 99.43 99.46 99.39 99.46
Atlas22 (D4) 70.52 73.77 70.15 73.67 74.22 75.35
CrossModa (D5) 99.85 99.76 99.72 99.78 99.74 99.81
Cosmos22 (D6) 72.60 96.47 94.48 96.89 96.95 97.45
ISLES22 (D7) 88.55 90.45 89.39 90.28 89.59 90.59
Hanseg (D8) 82.20 85.94 80.44 85.29 85.94 85.80
HNTS-MRG24 (D9) 71.83 71.26 73.47 67.88 73.22 74.07
BRATS24 Africa (D10) 95.66 95.36 94.95 94.94 95.33 95.06
T2 Aneurysms (D11) 62.24 55.56 51.79 58.97 59.38 61.18
Avg. NSD 82.04 84.08 82.39 84.17 84.95 85.58
Avg. NSD Rank 4.27 4.27 5.82 4.64 4.00 2.18
Figure 2: S3D ranks best across all methods. In addition to absolute mean performance, we report the ranking stability of the methods through bootstrapping for all test datasets as well as the aggregated rank across all datasets.
@@
Comparing the No (Dyn.) and No (Fixed) configurations, both trained from scratch, reveals that selecting the appropriate configuration for each dataset can significantly influence performance. For instance, on datasets D2 and D11, the dynamic configuration outperforms the fixed by +7 and +5 DSC points, respectively, while for D6, the fixed configuration yields results +18 DSC points higher. In the majority of datasets where the fixed configuration underperforms relative to the dynamic nnU-Net, pre-training helps to recover performance. However, in some cases, such as with D5, the dynamic default nnU-Net still proves superior.
@@ 5.2 Ablation experiments
Results are presented in Tab. 6. Our pre-trained S3D model leads to better downstream performance compared to training from scratch in this setting. With just 40 trained images, the fine-tuned model nearly matches the performance of the from-scratch model trained on the full dataset. Future research could explore whether optimizing the training duration or learning rate schedule could prevent the pre-trained network from overfitting during fine-tuning on such a limited number of training images.
@@
To assess the generalization capability of the proposed pre-training method, we tested two scenarios. First, we evaluated fine-tuning our method on an unseen modality using the TOF Angiography Aneurysms dataset (D12). As shown in Tab. 9 without pre-training, the fixed configuration suffers a performance drop of 20 Dice points. We attribute this to the significant difference in median spacing for the downstream task ($[0.50,0.43,0.43]$ mm), which has a higher resolution than the fixed target spacing of $[1,1,1]$ mm used in the pre-training experiments. This lower resolution likely increases the difficulty of segmenting small aneurysms. Despite this decrease, pre-training mitigates some of this degradation and proves highly beneficial compared to training the same configuration from scratch. Interestingly, ModelGenesis achieves the best results, potentially due to its use of intensity augmentations during pre-training, which increases robustness against brightness shifts, such as when generalizing to different MRI sequences. Second, we fine-tuned on the D2 dataset using only the T1 contrast-enhanced (T1ce) sequence and applied these models directly to D13 without any additional fine-tuning. While the dynamic configuration performed best on the in-distribution validation cases of D2, the results on D13 indicate that MAE pre-training improves generalization across different centers, with our S3D model yielding the best performance.
@@
Previous studies have demonstrated the positive impact of extended training schedules on the quality of learned representations for downstream tasks [14, 9]. To explore the relevance of this factor in the 3D medical domain, we conducted a similar experiment, evaluating training durations ranging from 62.5k to 1M steps. Our results indicate that the benefits of longer training schedules begin to degrade after 250k steps, as evident in Tab. 7.
@@
Initializing from pre-trained weights has the potential to reduce the computational resources needed for the network to adapt to new tasks. To assess this, we tested different fine-tuning durations on our development datasets. While maintaining 12.5k iterations for both the decoder and full network warm-up phases, we experimented with varying subsequent training lengths. As shown in Tab. 10, adding just 12.5k additional iterations (37.5k total) already outperforms training from scratch. However, achieving optimal performance still requires completing the full fine-tuning schedule.
@@
In many medical examinations, it is common to perform multiple scans, as clinicians often require images with different characteristics for accurate decision-making. Consequently, some datasets, like D2, D8, and D10, contain multiple input modalities. While pre-training may involve all modalities, we only feed one modality at a time into the network since not all patients have scans in every modality. This raises the question of how to handle datasets with multiple registered images. To address this, we conducted a 5-fold cross-validation on the D2 development dataset. We evaluated the replication of each input modality along with random initialization of the input stem weights. Additionally, we tested freezing the stem weights during the decoder warm-up phase. As shown in Tab. 8, the most stable and consistently effective approach was replicating the pre-trained stem and keeping it frozen during the decoder’s warm-up period.
@@ 6 Conclusion
This work is the first to demonstrate the potential of properly configured MAEs in 3D medical image segmentation. By overcoming key pitfalls in previous research, such as limited dataset sizes, outdated architectures, and insufficient evaluation, we show a consistent performance improvement over previous SSL methods. Notably, for the first time, we achieve consistent improvements over the dynamic, dataset-adaptive nnU-Net baseline, validated across a large and diverse set of development and testing datasets.
Limitations
In our current manuscript, we justify our selection of baseline methods by excluding prior work that was specific to architectures (3DMAE), methods that had no public code (GL-MAE), and methods that do not outperform a nnU-Net baseline trained from scratch (see Appendix Tab. 11).While our findings are promising, several avenues remain open for future exploration. We limited our method to Head and Neck MRI images, allowing us to focus on this region with a comparatively large pretraining dataset. Expanding to encompass the entire body across multiple imaging modalities would have required substantially more data to achieve comparable performance. Notably, in our ablations, we tested increasing training time and batch size (Appendix Tab. 12) which did not lead to performance improvements, but the question remains of whether scaling the pre-training dataset size or the model parameters could enable further improvements. Furthermore, the intensity shifts employed by Model Genesis SSL task hint at intriguing possibilities for improving generalization across unseen MRI modalities, which needs to be further explored for MAEs. Lastly, a data-centric approach to curating the most relevant data for SSL represents an exciting frontier for future research. Raw clinical datasets often contain images not intended for diagnostic purposes, such as those used for scanner calibration, which can dilute the effectiveness of pre-training. While we applied basic filtering to exclude low-quality data, more sophisticated filtering techniques could significantly enhance the quality of the pre-training process.
To conclude, this work follows the spirit of prior studies like nnU-Net [18] showing that a robust development strategy, informed model configuration, and rigorous validation leads to true and sustainable performance improvements, contrasting the current hype for employing and modifying the latest network architecture. With this paper, we hope to contribute to a cultural shift in the community toward validation-driven development enabling true scientific progress.
@@ Appendix A Method configuration
Across all baseline methods, we utilize a common set of hyperparameters. For all baseline methods we utilize the same pre-training dataset with the same image preprocessing. Moreover we use the same amount of pre-training steps (250k) as for our S3D-B method and the same fine-tuning scheme, as highlighted in Tab. 10. Aside from this, we employ the SGD optimizer with LR 1e-2 with a PolyLR schedule, momentum 0.99 and weight decay 3e-5 across all pre-training experiments, as they showed to be highly robust and reliable in the supervised medical image segmentation setting using CNNs [18]. Moreover, we denote that all these methods have their backbones replaced with a ResEnc U-Net to minimize confounding effects of different architectures.
@@ A.1 Models Genesis
Models genesis [44] pre-text task is centered around restoring original patches from transformed versions. The transformed version is achieved by applying four different transformations in various combinations, with the following transformations: a composition of four separate pre-training schemes: (i) Non-linear intensity transformation:Alters the intensity distribution while preserving the anatomy, focusing on learning the appearance of organs. (ii) Out-painting:Removes part of the image and requires the model to extrapolate from the remaining image, forcing it to learn the global structure of the organs. (iii) In-painting:Masks a part of the image, and the model learns to restore the missing parts, focusing on local continuity and context. After having transformed the original image through these augmentations, the model is trained to recover the original image through a convolutional encoder-decoder architecture. This approach consolidates different tasks (appearance, texture, and context learning) into one unified image restoration task, making the model more robust and generalizable.
@@
The entire set of hyperparameters of Models Genesis are contained within the data-augmentation. This allows us to transfer this transformation pipeline, as provided in the official repository without any changes to the hyperparameters.